HIGH RESOLUTION OVERTONE SPECTROSCOPY OF THE ACETYLENE VAN DER WAALS DIMER, 12(C_{2}H_{2})_{2}

\underline{K. DIDRICHE}, C. LAUZIN, T. FOLDES, D. GOLEBIOWSKI, M. HERMAN, Service de Chimie quantique et Photophysique CP160/09, Faculté des Sciences, Université Libre de Bruxelles (U.L.B.), Av. Roosevelt, 50, B-1050, Bruxelles, Belgium; C. LEFORESTIER, ACTMM-CC 15.01, Institut Charles Gerhardt, 34095 Montpellier, France.

CW-cavity ring down spectroscopy was used to record in a free jet expansion the spectrum of the absorption band in (^{12}C_{2}H_{2})_{2} with origin at 6547.6 cm^{−1}. It is a perpendicular band and corresponds to 2CH excitation in the hat unit of the T-shaped dimer. Calibration (better than ±1 × 10^{−3} cm^{−1} accuracy) and ring-down time (130 μs) were improved compared to a previous contribution ^{a}, and a line-by-line analysis was undertaken. Three series of lines were identified involving levels with A1^{+}, E^{+} and B1^{+} tunneling symmetries, confirming the spectral and symmetry analyses reported in the literature for the 1CH excitation band ^{b}. 164 vibration-rotation-tunneling lines were assigned in the K′_{a} − K′′_{a} = 2 − 3, 0 − 1, 2 − 1 and 4 − 3 sub-bands and effective rigid rotor vibration-rotation constants were obtained by simultaneously fitting 1CH and 2CH lines from the same symmetry series. Perturbations affecting the K_{a} stacks, in particular, are reported. The tunneling frequency in 2CH is estimated to be ν_{tunn.}^{2CH} = 270 MHz for the K_{a} = 0 stack. The rotational temperature is determined to be 23 K from relative line intensities and the half lifetime of the dimer in the 2CH hat state is estimated to 1 ns from individual line widths.

aDidriche et al., Mol. Phys. 2010, 108, 2158-2164

bFraser et al., J. Chem. Phys. 1988, 89, 6028-6045